# LLM-Enhanced Multi-Agent Reinforcement Learning with Expert Workflow for Real-Time P2P Energy Trading

### v2

@@ Abstract
Real-time peer-to-peer (P2P) electricity markets dynamically adapt to fluctuations in renewable energy and variations in demand, maximizing economic benefits through instantaneous price responses while enhancing grid flexibility. However, scaling expert guidance for massive personalized prosumers poses critical challenges, including diverse decision-making demands and lack of customized modeling frameworks. This paper proposed an integrated large language model-multi-agent reinforcement learning (LLM-MARL) framework for real-time P2P energy trading to address challenges such as the limited technical capability of prosumers, the lack of expert experience, and security issues of distribution networks. LLMs are introduced as experts to generate personalized strategy, guiding MARL under the centralized training with decentralized execution (CTDE) paradigm through imitation learning. A differential attention-based critic network is designed to enhance convergence performance. Experimental results demonstrate that LLM generated strategies effectively substitute human experts. The proposed multi-agent imitation learning algorithms achieve significantly lower economic costs and voltage violation rates on test sets compared to baselines algorithms, while maintaining robust stability. This work provides an effective solution for real-time P2P electricity market decision-making by bridging expert knowledge with agent learning.
@@ I INTRODUCTION
The rise of peer-to-peer (P2P) energy trading has shifted electricity users from traditional consumers to ”prosumers,” combining both production and consumption [1]. However, this development faces two challenges: on the virtual layer, prosumers often lack the technical capability for repeated trading and efficient energy management [2]; on the physical level, ensuring system security during the transmission of electricity transactions from the virtual layer in actual distribution networks remains a challenge [3]. Additionally, traditional model-driven approaches fail to meet the dynamic response requirements of real-time P2P trading, resulting in a conflict between computational efficiency and decision-making flexibility [4]. Thus, reinforcement learning (RL) has become a promising solution.
RL, due to its dependence on environmental interactions, faces challenges due to slow convergence and difficulty of reaching a global optimal solution [5]. Additionally, poor decision-making in the early training stages can endanger the security of the power grid. In response to these issues, mixed-integer programming-based expert solvers and imitation learning methods have improved the training efficiency of microgrid operation [6] and distribution network reconfiguration [7]. However, the decision-making driven by a single expert’s experience has limitations in adapting to the personality of prosumers in the P2P energy trading, as standardized strategies struggle to meet their diverse needs. Moreover, relying on a human Distribution System Operator (DSO) as the decision-making entity not only encounters practical limitations such as high labor costs and low response efficiency, but its inherent lack of generalization further exacerbates operational risks [8].
@@
With the emergence of large language models (LLMs) like ChatGPT, LLMs showcase strong reasoning, decision-making, and generalization abilities, addressing the shortcomings of using single experts in RL to handle the heterogeneity of prosumers, as well as the challenges of human expert labor costs and generalization. While LLMs have begun to be applied in the power and energy sector—such as in [9], where LLMs replace humans for designing penalty functions in safe RL, [10] evaluates LLMs’ performance in various power system tasks, highlighting their potential in complex system modeling and reasoning. The study in [11] introduces a power multi-agent framework with a feedback mechanism, validated on the DALINE and MATPOWER platforms. However, the integration of LLMs as expert systems for assisting prosumers in RL training for P2P energy trading remains underexplored. Currently, in non-power system domains LLMs have been successfully used as expert guides in autonomous driving RL training [12, 13], providing valuable insights for the application of LLMs in the P2P energy trading domain.
@@
Multi-agent Reinforcement Learning (MARL) stands out among RL approaches for P2P energy trading due to its capability to model strategic interactions among distributed prosumers in continuous action spaces, enabling decentralized decision-making and fostering collaborative behaviors. However, the complexity of managing numerous agents with high-dimensional actions under the Centralized Training and Decentralized Execution (CTDE) framework[14] limits the ability to exploit global collaborative information [15]. To improve learning efficiency, attention mechanisms have been introduced into MARL models [16] to better extract relevant features. Empirical studies confirm their effectiveness in power system applications such as voltage regulation [17], microgrid trading [18], and community-based P2P trading [19]. Nevertheless, conventional attention mechanisms often lacks a holistic understanding of the entire sequence and may ignore critical information [20], which constrains MARL’s performance in real-world P2P trading—an issue this work seeks to address.
In summary, this paper proposes an innovative integration framework that guides MARL training through LLM experts within the context of maximizing social welfare in local P2P real-time electricity markets involving prosumer collaboration. At the initial stage of training, the LLM generates model code based on differentiated prosumers’ demands. Each prosumer’s expert is embedded within a multi-LLM workflow, where real-time states are passed through to commercial solvers, indirectly generating expert strategies for each prosumer.
During the training phase, a CTDE-based MARL imitation learning algorithm is proposed. Moreover, inspired by the recent application of Differential Transformer in LLMs [21], we design an enhanced critic network architecture based on the differential attention. This architecture is specifically developed to mitigate irrelevant information interference from other agents during training and improve the overall convergence performance of the algorithm. The main contributions of this article are listed as follows:
@@
    A novel LLM-MARL integrated framework is proposed for the real-time P2P electricity market. By introducing LLMs as expert in the P2P energy trading, the framework replaces human experts to guide MARL agents during training. This significantly reduces manual intervention costs and achieves a deep integration of expert knowledge and LLM-based reasoning.
@@
    An LLM expert workflow tailored to local P2P electricity markets is developed for each prosumer. The workflow includes model generation, tool retrieval, code generation, and code correction. By processing state information, it dynamically generates prosumer strategies that balance economic performance and distribution network security, thus providing reliable expert guidance during training.
@@
    A novel multi-agent imitation learning algorithm is proposed. It introduces the Wasserstein metric to measure the similarity between expert strategy and agent policy, enabling effective guidance from the LLM experts’ workflow. Furthermore, a differential multi-head attention-based Critic network is designed to improve policy evaluation accuracy and accelerate the learning process, thereby boosting overall algorithmic performance.
@@ II-A Prosumer energy management and P2P Trading
Prosumers are modeled as independent energy units with conventional distributed generators (CDGs), renewable distributed generators (RDGs), battery energy storage systems (BESSs), and controllable loads (CLs). A 15-minute interval is adopted for optimization, aligning with real-time P2P electricity market practices.
@@
$\lambda_{t}^{P2P}=\lambda^{P2P}(\lambda_{t}^{B}-\lambda_{t}^{S})+\lambda_{t}^{S},$ (24)
where $\lambda^{P2P}\in[0,1]$ is a price coefficient that determines the fairness of the P2P market. This pricing scheme incentivizes local trading by offering sellers a higher price than grid export rates and buyers a lower price than grid retail rates.
@@ II-B Dec-POMDP for P2P Energy Trading
In P2P energy trading, the inherent uncertainty challenges traditional mathematical optimization methods in meeting the precision and real-time demands of prosumer optimization control. RL methods can address these limitations, enabling data-driven optimization decisions. This paper models the P2P trading problem for multiple prosumers in a distribution network as a Dec-POMDP, represented as an eight-tuple $\langle I,A,S,O,P,r,\pi,\gamma\rangle$, where $S$ is the global state space, $A$ is the joint action space, $r$ is the global reward function based on state transitionsPP, $O$ is the observation space, and $\gamma$ is the discount factor. The model captures the features of information asymmetry and decentralized decision-making through partial observability and decentralized architecture.
@@
    State $S$: The global state at time $t$, denoted as $s_{t}=\{s_{i,t}\,|i\in{N}\}$, for $\forall s_{t}\in{N}$, encompasses the operational conditions of all nodes in the distribution network, Specifically, $s_{i,t}$ includes the operational state of the prosumer, the distribution network’s interaction state, and the previous action of the prosumer:
@@
    All prosumers aim to achieve social welfare maximization by avoiding distribution networks voltage violations and minimizing costs. To this end, a unified global reward function is shared among all agents:
@@ III METHODOLOGY
This paper focuses on P2P energy trading in distribution networks and introduces a novel MARL framework constrained by expert strategies. The proposed framework adopts an off-policy RL paradigm to enable efficient online updates. In this approach, LLM serves as an expert, generating personalized strategies to guide prosumers in energy trading. Fundamentally, it constitutes a form of “phase-triggered code synthesis training” enabled by LLM. Specifically, at the onset of each training phase, the LLM synthesizes optimization code tailored to the current system configuration. During subsequent online learning iterations, this code is executed in a loop to produce state-dependent, personalized strategies for each prosumer agent. To ensure a principled integration of expert knowledge with agent learning, the MARL algorithm is enhanced using the Wasserstein metric, which promotes alignment between the LLM strategy and the learned agent policies. When prosumer devices are modified or newly added, the LLM regenerates the code to adapt to the updated environment. As a result, the trained agent is capable of executing energy trading tasks autonomously without further LLM intervention. The overall framework is illustrated in Figure 1.
@@ III-A1 Knowledge Enhancement Methods
Knowledge injection enhance the capabilities of LLMs and help mitigate information hallucination when these models act as domain experts. The main knowledge enhancement approaches include Supervised Fine-Tuning (SFT) and Retrieval-Augmented Generation (RAG). Unlike SFT, which requires full parameter fine-tuning, RAG decouples the knowledge storage from the inference generation. It only requires maintaining an external knowledge base. This characteristic not only ensures the real-time model outputs but also improves the interpretability of the generated results through explicit knowledge tracing. Recent research from Microsoft has demonstrated that RAG outperforms SFT in integrating domain-specific knowledge into LLMs [24], making RAG the core knowledge enhancement framework in this paper.
@@ III-A2 Prosumer-Centric LLM Expert Strategies
This paper presents an LLM-based expert execution workflow to support personalized operational strategy for each prosumer. The system includes four complementary LLM agent expert and a module for distribution networks security verification via DSO, All optimization models are formulated in cvxpy and solved using commercial solvers such as Gurobi [27], CPLEX [28], or MOSEK [29]. The overall process is shown in the figure2, as described below:
@@
  • Code Generation Expert: This module LLM constructs the optimization model in the cvxpy framework using knowledge of the power domain, the cvxpy programming syntax, and the device parameters and state data of the prosumer. It outputs a DCP rules correct and feasible modeling code.
@@ III-B Multi-Agent Imitation Learning Algorithm
For the prosumer collaborative learning problem, this paper proposes an expert strategy-constrained multi-agent imitation learning Algorithm based on the CTDE framework. The Algorithm constructs a joint state-action value function (Q-function) and state value function (V-function) through a centralized evaluation network, integrating the global environmental state and the behavior policies of all agents during the training phase, thereby guiding the differential optimization of individual agent policies. During execution, a decentralized approach is adopted, where each agent independently makes decisions based on local observations through independently trained networks, balancing cooperative benefits and decision-making efficiency.

### v3

@@ Abstract
Real-time peer-to-peer (P2P) electricity markets dynamically adapt to fluctuations in renewable energy and variations in demand, maximizing economic benefits through instantaneous price responses while enhancing grid flexibility. However, scaling expert guidance for massive personalized prosumers poses critical challenges, including diverse decision-making demands and a lack of customized modeling frameworks. This paper proposes an integrated large language model-multi-agent reinforcement learning (LLM-MARL) framework for real-time P2P energy trading to address challenges such as the limited technical capability of prosumers, the lack of expert experience, and security issues of distribution networks. LLMs are introduced as experts to generate personalized strategies, guiding MARL under the centralized training with decentralized execution (CTDE) paradigm through imitation. To handle the scalability issues inherent in large-scale P2P networks, a differential attention-based critic network is introduced to efficiently extract key interaction features and enhance convergence. Experimental results demonstrate that LLM-generated strategies effectively substitute human experts. The proposed imitative expert MARL algorithms achieve significantly lower economic costs and voltage violation rates on test sets compared to baseline algorithms, while maintaining robust stability. This paper provides an effective solution for the real-time decision-making of the P2P electricity market by bridging expert knowledge with agent learning.
@@ I INTRODUCTION
The rise of peer-to-peer (P2P) energy trading has shifted electricity users from traditional consumers to ”prosumers,” combining both production and consumption [1]. However, this development faces two challenges: on the virtual layer, prosumers often lack the technical capability for repeated trading and efficient energy management [2]; on the physical level, ensuring system security during the transmission of electricity transactions from the virtual layer in actual distribution networks remains a challenge [3]. Although the majority of energy trading is settled in the day-ahead stage, actual load and renewable generation often deviate substantially from the scheduled profiles due to inherent uncertainties. Consequently, real-time mechanisms are indispensable for rapidly rescheduling trades to restore power balance and enabling distributed resources to provide immediate balancing services[4]. At present, electricity markets such as Nord Pool in the Nordic region, PJM and ERCOT in the United States, and Australia’s National Electricity Market mainly operate under a two-stage framework with separate settlement in the day-ahead and real-time markets. As the real-time market is a critical component of electricity market operations, it is necessary to operate P2P market mechanisms in a real-time operational mode as well[5].
In response to this requirement, recent studies have focused on developing real-time P2P energy trading. Specifically, studies such as [6] and [7] employed Lyapunov optimization to relax temporal coupling constraints and utilized the Alternating Direction Method of Multipliers to achieve distributed solutions for real-time P2P energy rescheduling. In parallel, [8] proposed a real-time P2P energy trading method grounded in Model Predictive Control (MPC). Nevertheless, these approaches have inherent limitations. Lyapunov optimization inevitably introduces approximation errors due to the relaxation of temporal coupling constraints. Similarly, the open-loop nature of MPC policies renders them intrinsically suboptimal in the presence of stochastic noise [9]. Furthermore, such conventional model-driven methods rely heavily on precise system modeling and parameterization, which are typically difficult to obtain in practice [10]. Consequently, they often struggle to accommodate the highly dynamic nature of real-time P2P energy trading, creating a trade-off between computational efficiency and decision-making flexibility [11]. To address these challenges, reinforcement learning (RL) has emerged as a promising solution, owing to its robust adaptability to uncertain environments and its ability to facilitate rapid strategy generation [12].
RL relies heavily on trial-and-error interactions with the environment, faces fundamental challenges when applied to P2P energy trading systems with strict operational constraints. RL methods struggle to efficiently converge to a global market equilibrium [13], particularly in large-scale P2P markets characterized by highly personalized prosumers, whose preferences and capacities demand individualized and incentive-compatible decision-making rather than uniform policies [14]. To mitigate these challenges, optimization-based solvers and imitation learning techniques have been introduced to incorporate expert knowledge into the learning process, thereby improving training efficiency and solution feasibility in related power system applications, such as microgrid operation [15] and distribution network reconfiguration [16]. For instance, [17] proposed a multi-virtual-agent imitation learning framework that leverages adversarial imitation across multiple virtual environments to achieve robust microgrid energy scheduling under uncertain power supply interruptions. However, imitation learning approaches driven by a single expert’s decision-making experience remain limited in their ability to adapt to the diverse personalities and strategic preferences of prosumers in P2P energy trading. Standardized expert strategies often fail to capture individualized behaviors and incentive structures. Moreover, relying on a human Distribution System Operator (DSO) as the expert decision-maker introduces practical constraints, including high labor costs and limited real-time responsiveness, while the lack of generalization capability further amplifies operational risks [18].
@@
With the emergence of large language models (LLMs) like ChatGPT, LLMs showcase strong reasoning, decision-making, and generalization abilities, addressing the shortcomings of using single experts in RL to handle the heterogeneity of prosumers, as well as the challenges of human expert labor costs and generalization. While LLMs have begun to be applied in the power and energy sector—such as in [19], where an $RL^{2}$ mechanism is designed with LLM-assisted safe reinforcement learning for active distribution network energy management via an iterative penalty function, and in [20], where the LLM is used to interpret linguistic operational stipulations and generate reward signals, alleviating the need for manually designing explicit reward functions for qualitative objectives. [21] evaluates LLMs’ performance in various power system tasks, highlighting their potential in complex system modeling and reasoning. The study in [22] introduces a power multi-agent framework with a feedback mechanism, validated on the DALINE and MATPOWER platforms. However, the integration of LLMs as expert systems for assisting prosumers in RL training for P2P energy trading remains underexplored. Currently, in non-power system domains LLMs have been successfully used as expert guides in autonomous driving RL training [23, 24], providing valuable insights for the application of LLMs in the P2P energy trading domain.
Nevertheless, applying single-agent RL to P2P energy trading entails inherent limitations. When consumers seek individual economic optimality, uncoordinated decisions can cause an excessive concentration of load or generation, posing risks to the security of the distribution network [25]. Furthermore, the concurrent and independent policy updates of multiple prosumers render the learning environment highly non-stationary from the perspective of any single agent [26]. To address these challenges, multi-agent reinforcement learning (MARL) has emerged as a promising paradigm for P2P energy trading, as it explicitly captures strategic interactions among distributed prosumers in continuous action spaces, supports decentralized decision-making, and promotes cooperative behaviors. [27] proposed a multi-agent adversarial reinforcement learning to solve the active voltage control problem in peer-to-peer energy trading-enabled distribution networks. [28] formulates the network-constrained MARL P2P energy trading problem as a cooperative Markov game.
@@
However, the complexity of managing numerous agents with high-dimensional actions under the Centralized Training and Decentralized Execution (CTDE) framework[29] limits the ability to exploit global collaborative information [30]. To improve learning efficiency, attention mechanisms have been introduced into MARL models [31] to better extract relevant features. Empirical studies confirm their effectiveness in power system applications such as voltage regulation [32], microgrid trading [33], and community-based P2P trading [34]. While attention is inherently a neural network architecture rather than a market mechanism, it is particularly well-suited for modeling the uneven interaction patterns in P2P energy trading[35]. Nevertheless, conventional attention mechanisms often lacks a holistic understanding of the entire sequence and may ignore critical information [36]. In a distribution network, a prosumer’s state is significantly affected only by a subset of relevant peers, rather than the entire population. Standard attention mechanisms may fail to distinguish these critical trading partners from irrelevant ones, leading to inefficient coordination in large-scale P2P systems, which constrains MARL’s performance in real-world P2P trading.
In summary, this paper addresses the P2P energy trading limitations of existing reinforcement learning methods identified in TABLE I. It proposes a novel integrated framework in which LLM-based experts guide the training of MARL agents, with the goal of maximizing social welfare in local real-time P2P electricity markets featuring prosumer collaboration. During the initial stage of training, the LLM generates model code tailored to differentiated prosumers’ demands. Specifically, each prosumer’s expert is embedded within a multi-LLM workflow, where real-time states are fed into solvers, indirectly generating expert strategies for each prosumer. For the training phase, a CTDE-based imitative expert MARL algorithm is proposed. Furthermore, inspired by the recent application of the Differential Transformer in LLMs [37], an enhanced critic network architecture utilizing differential attention is designed. This architecture is specifically developed to mitigate irrelevant information interference from other agents among large-scale prosumers during training and to enhance the overall convergence performance of the algorithm. The main contributions of this article are summarized as follows:
TABLE I: Comparison limitations in P2P energy trading reference
Ref. Prosumer Personalized Network Constraints Large Scalability Method
[14] $\checkmark$ $\times$ RL
[28] $\times$ $\checkmark$ $\times$ MARL
[34] $\times$ $\times$ $\checkmark$ MARL
[35] $\times$ $\times$ $\checkmark$ MARL
This Paper $\checkmark$ $\checkmark$ $\checkmark$ LLM-MARL
@@
    A novel LLM-MARL integrated framework is proposed for the real-time P2P electricity market. By introducing LLMs as expert in the P2P energy trading, the framework replaces human experts to guide MARL agents during training. This directly substitutes the process of human expert code generation or expert guidance and achieves a deep integration of expert knowledge and LLM-based reasoning.
@@
    An LLM expert workflow tailored to local P2P energy trading is developed for each prosumer. This workflow transforms prosumer natural language input into executable actions through model generation, tool retrieval, code generation, and code correction. By processing state information, it dynamically generates prosumer strategies that balance economic performance and distribution network security, thereby providing reliable expert guidance during training.
@@
    A novel imitative expert MARL algorithm is proposed. It introduces the Wasserstein metric to measure the similarity between expert strategy and agent policy, enabling effective guidance from the LLM experts’ workflow. Furthermore, a differential multi-head attention-based Critic network is designed to improve policy evaluation accuracy and accelerate the learning process in the large-scale P2P energy trading, thereby boosting overall algorithmic performance.
@@ II-A Prosumer energy management and P2P Trading
Prosumers are modeled as independent energy units with conventional distributed generators (CDGs), renewable distributed generators (RDGs), battery energy storage systems (BESSs), and controllable loads (CLs). This study adopts a 15-minute interval for optimization to align with real-time P2P energy trading settlement practices[6].
@@
$\lambda_{t}^{P2P}=\kappa^{P2P}(\lambda_{t}^{B}-\lambda_{t}^{S})+\lambda_{t}^{S},$ (24)
where $\kappa^{P2P}\in[0,1]$ is a price coefficient that determines the fairness of the P2P market. This pricing scheme incentivizes local trading by offering sellers a higher price than grid export rates and buyers a lower price than grid retail rates.
@@ II-B Dec-POMDP for P2P Energy Trading
In P2P energy trading, the inherent uncertainty challenges traditional mathematical optimization methods in meeting the precision and real-time demands of prosumer optimization control. RL methods can address these limitations, enabling data-driven optimization decisions. This paper models the P2P trading problem for multiple prosumers in a distribution network as a Dec-POMDP, represented as an eight-tuple $\langle I,A,S,O,P,r,\pi,\gamma\rangle$, where $S$ is the global state space, $A$ is the joint action space, $r$ is the global reward function based on state transitions $P$, $O$ is the observation space, and $\gamma$ is the discount factor. The model captures the features of information asymmetry and decentralized decision-making through partial observability and decentralized architecture.
@@
    State $S$: The global state at time $t$, denoted as $s_{t}=\{s_{i,t}\,|i\in{N}\}$, for $\forall s_{t}\in{N}$, encompasses the operational conditions of all nodes in the distribution network. $V_{i,t-1}$ captures node voltage magnitudes via phasor measurement units and wireless sensors[40], enabling realistic prosumer state observation. Specifically, $s_{i,t}$ includes the operational state of the prosumer, the distribution network’s interaction state, and the previous action of the prosumer:
@@
    Given the strong coupling between active power injections and nodal voltages in P2P energy trading, ignoring physical limits renders market outcomes infeasible[27]. Following [28], we adopt a unified global reward function shared among all agents. This design is crucial because, in physically coupled networks, a shared signal is necessary to guide agents toward a Nash Equilibrium that maximizes social welfare while strictly adhering to voltage safety limits. To this end, the reward function is designed as:
@@ III METHODOLOGY
This paper focuses on P2P energy trading in distribution networks and introduces a novel MARL framework constrained by expert strategies. The proposed framework adopts an off-policy RL paradigm to enable efficient online updates. In this approach, LLM serves as an expert, generating personalized strategies to guide prosumers in energy trading. Fundamentally, it constitutes a form of “phase-triggered code synthesis training” enabled by LLM. Specifically, at the onset of each training phase, the LLM synthesizes optimization code tailored to the current system configuration. During subsequent online learning iterations, this code is executed in a loop to produce state-dependent, personalized strategies for each prosumer agent. To ensure a principled integration of expert knowledge with agent learning, the imitative expert MARL algorithm is enhanced using the Wasserstein metric, which promotes alignment between the LLM strategy and the learned agent policies. When prosumer devices are modified or newly added, the LLM regenerates the code to adapt to the updated environment. The overall framework is illustrated in Fig. 1.
@@ III-A1 Knowledge Enhancement Methods
Knowledge injection enhances the capabilities of LLMs and help mitigate information hallucination when these models act as domain experts. The main knowledge enhancement approaches include Supervised Fine-Tuning (SFT) and Retrieval-Augmented Generation (RAG). Unlike SFT, which requires full parameter fine-tuning, RAG decouples the knowledge storage from the inference generation. It only requires maintaining an external knowledge base. This characteristic not only ensures the real-time model outputs but also improves the interpretability of the generated results through explicit knowledge tracing. Recent research from Microsoft has demonstrated that RAG outperforms SFT in integrating domain-specific knowledge into LLMs [41], making RAG the core knowledge enhancement framework in this paper.
@@ III-A2 Prosumer-Centric LLM Expert Strategies
This paper presents an LLM-based expert execution workflow to support personalized operational strategy for each prosumer. The system includes four complementary LLM agent experts and a module for distribution networks security verification via DSO, All optimization models are formulated in cvxpy and solved using mathematical optimization solvers such as Gurobi [44]. The overall process is shown in Fig. 2, as described below:
@@
  • Code Generation Expert: This module LLM constructs the optimization model in the cvxpy framework using knowledge of the power domain, the cvxpy programming syntax, and the device parameters and state data of the prosumer. It outputs DCP-rule-compliant and feasible modeling code.
@@ III-B Multi-Agent Imitation Learning Algorithm
For the prosumer collaborative learning problem, this paper proposes an expert strategy-constrained MARL Algorithm based on the CTDE framework. The Algorithm constructs a joint state-action value function (Q-function) and state value function (V-function) through a centralized evaluation network, integrating the global environmental state and the behavior policies of all agents during the training phase, thereby guiding the differential optimization of individual agent policies. During execution, a decentralized approach is adopted, where each agent independently makes decisions based on local observations through independently trained networks, balancing cooperative benefits and decision-making efficiency.